one half of a certain number is equal to 65% of the second number. find the ratio of 1st and 2nd number?

Answer by LinnW(1048)

set f = the first number
s = the second number
(1/2)f = 0.65s
multiply each side by 2
f = 1.30s
f/s = 1.30s/s = 1.30 = 130/100 = 13/10